# PixelArena: A benchmark for Pixel-Precision Visual Intelligence

### v1

@@ 1 Color Palettes and Prompts
We also present the standard color palette used in experiments on the subset of COCO [coco-dataset] in Fig. 13. We intentionally split the color palette into seven images so that none of them is excessively tall, as such a tall image may be outside a model’s training distribution or may break its vision module. The prompt (Listing. LABEL:coco-prompt) for the COCO experiments is similar to the one in Listing. LABEL:prompt with a few minor modifications.
Figure 13: Seven Images of the standard color palette of the COCO dataset.
Listing 2: Prompt Template for our COCO experiments. We omit the rest of color codings here.
I want you to do semantic segmentation based on the given category labels.
The label encodings are
other : [0, 0, 0]
... omitted
For your convenience, Ive also given you a color palette (the rest of the images) for the label encodings.
You can first recognize all categories of all subjects in the first image and then draw the mask. Note that the first category other is used only when therere no suitable category labels for a subject in the image.
2 Sampling Parameters
We report the sampling parameters of MLLMs used in our experiments in Table. 1. We did not purposefully tweak the relevant sampling parameters; rather, we used the defaults in the APIs or source code.
Table 1: Sampling Parameters for MLLMs
MLLM Temperature Top P
gmn3 1.0 0.95
gmn25 1.0 0.95
gpti NA NA
emu35 1.0 1.0
unimoe2-omni 1.0 1.0
unimoe2-image 1.0 1.0
3 More quantitative results
@@
Figure 14: mIoU scores for experiments on celeb. For $p=[3,5]$, we ask MLLMs to try 3 or 5 times and select the best result in these attempts. As sam3 and segface contain no stochastic components, we did not run more attempts. Due to their poor performance, we did not run experiments of emu35 and unimoe2 with more attempts.
Figure 15: Dice Scores for experiments on celeb. For $p=[3,5]$, we ask MLLMs to try 3 or 5 times and select the best result in these attempts. As sam3 and segface contain no stochastic components, we did not run more attempts. Due to their poor performance, we did not run experiments of emu35 and unimoe2 with more attempts.
@@
Figure 16: mIoU Scores for experiments on coco. For $p=[3,5]$, we ask MLLMs to try 3 or 5 times and select the best result in these attempts. As 1former contain no stochastic components, we did not run more attempts. Due to their poor performance in previous experiments, we did not run experiments for other MLLMs with more attempts.
Figure 17: Dice Scores for experiments on coco. For $p=[3,5]$, we ask MLLMs to try 3 or 5 times and select the best result in these attempts. As 1former contain no stochastic components, we did not run more attempts. Due to their poor performance in previous experiments, we did not run experiments for other MLLMs with more attempts.

### v2

@@ 3.2 Models and Mask Generation
Listing 1: Prompt Template for our celeb experiments. We omit the rest of color codings here. As there is ambiguity in the left and right in terms of references (i.e., with respect to images or persons), we clarify this in length in the prompt to avoid confusion.
For OMMs: We select recent models with strong image generation capabilities, including Gemini 3 Pro Image (gmn3) DeepMind (2025b), Gemini 2.5 Flash Image (gmn25) DeepMind (2025a), GPT Image 1 (gpti) OpenAI and et al. (2024), Emu 3.5 (emu35) Cui et al. (2025), and Uni-MoE-2 (unimoe2) Li et al. (2025). Note that we tested two variants of unimoe2: Uni-MoE-2 Omni (unimoe2-omni), the flagship model of the series, and Uni-MoE-2 Image (unimoe2-image), the variant finetuned for image generation. As they natively generate images instead of label vectors, we first prompt them to generate images with specified color encodings and then convert the pixels from RGB values to segmentation class labels. The prompts for the two datasets are composed of three parts: an image from the dataset, an image of the color palette of label encodings (Fig. 1 and Fig. 2 in Suppl. A ), and a short text (Listing 1). These prompts provide task specifications and visual grounding for the color encodings, as well as some clarifications. Note that no examples are given in our prompts, meaning the models have to learn SS tasks zero shot. For the sampling parameters of the OMMs, refer to Table 1 in Suppl. B. Given the mask images generated by OMMs, we compare the RGB value of each pixel with the color encodings for the labels, selecting the nearest color and label. It is formulated as Eq. 1, where $\vec{e_{i}}$ is the RGB color vector of a label with index $i$, and $\vec{p}$ is the color vector of a pixel.
$i=\operatorname*{argmin}_{i}\ (\vec{e_{i}}\cdot\vec{p})$ (1)
For SAM 3: SAM 3 (sam3) Carion et al. (2025) accepts text as the prompt for mask generation. We prompt sam3 with the labels of CelebAMask-HQ Lee et al. (2020) one by one and merge the corresponding 19 masks into one final mask. For the label of each pixel in the overlapping areas of these masks, we randomly pick one from the overlapping labels.
For specialized computer vision models: We use the pretrained ConvNext Liu et al. (2022) variant of SegFace (segface) Narayan et al. (2025) as a strong baseline model on celeb; while on coco, we use OneFormer (1former) Jain et al. (2023).
4 Analysis
Due to the stochastic components (e.g., token sampling, diffusion module) in OMMs, the generation of mask images is inherently stochastic. Therefore, we present results from multiple attempts. The number of attempts is $p=[1,3,5]$.
4.1 Qualitative Comparisons
Figure 2: Comparison between the Results of Different Models on celeb. The short code names are shown on the top of images. The results are not cherry-picked.
Figure 3: Best prediction across celeb by gmn3 with F1 score $0.708$. The short code names are shown on the top of images.
Figure 4: Worst prediction across celeb by gmn3 with F1 score $0.081$ and parallel attempts. Attempt numbers are shown on the top of images.
In Fig. 2, we present the results of different models on celeb for qualitative comparison. Among all OMMs, gmn3 is the only one that understands the task requirements and completes it with high quality. gpti and gmn25 partially understand the task, but gmn25 lacks precise color control or fails to understand the color encodings, while gpti lacks precise control over the composition of the image and hallucinates the upper body of the person. As for sam3, it sometimes misses some labels. emu35 and unimoe2 models completely misunderstand the task while presenting different failure patterns. emu35 failed to draw plausible masks, but it could control its image generation process to replicate most features of the original image. In contrast, unimoe2 models could not even draw an image similar to the original image, which may be due to its vision system failing to capture the original image, failing to propagate the visual information to its generation module, or failing to control its generation process.
Such failure modes are common in the results of the respective models, implying a potential misalignment between the vision system and the generation module, or a lack of control over the generation process.
We further investigate the results of gmn3. We present the best and worst results from the model in Fig. 3 and Fig. 4. The best prediction has a nearly indistinguishable difference from the reference mask, while the worst prediction is dramatically low in quality. However, the other four attempts in Fig. 4 present reasonable results, which suggests that the generation process is not stable or robust.
Figure 5: Comparison between the reference mask and masks predicted by two strong models on celeb. Their short code names and F1 scores are on the top of the images. We picked the image on which segface achieved the highest F1 score $0.9539$ while gmn3 achieved $0.6906$.
Although we do not intend to use OMMs to compete with the specialized computer vision model (e.g., segface), we present the result in which segface achieved the highest F1 score $0.9539$, while gmn3 achieved $0.6906$ in Fig. 5. We find that these two masks are visually very similar, while the scores have a significant gap.
4.2 Quantitative Results and Examining Data Contamination
@@
Figure 6: F1 Scores for experiments on celeb. For $p=[3,5]$, we ask OMMs to try 3 or 5 times and select the best result in these attempts. As sam3 and segface contain no stochastic components, we did not run more attempts. Due to their poor performance, we did not run experiments of emu35 and unimoe2 with more attempts.
We present the F1 scores in Fig. 6. For mIoU and Dice, please refer to Figs. 3 and 4 in Suppl. C. Aligned with our qualitative analysis, gmn3 achieved the best F1 score, mIoU, and Dice in the selected OMMs, although it lags behind segface.
As the results of gmn3 are surprisingly good on celeb, we further check whether data contamination is the cause of such good performance rather than true generalization, since CelebAMask-HQ Lee et al. (2020) has published all images and masks on the Internet. We shuffle the color encodings (Fig. 1 in Suppl. A) instead of using the standard encodings in a new experiment. Noteworthy is that, as shown in Fig. 6, after we shuffled the color encodings, the performance of gmn3 (gmn3-shuffled) did not drop but instead increased by roughly 10% compared to its original result. This means the model did not memorize the reference masks but truly understood the task, including the required color mapping using arbitrary color encodings.
4.3 Further Failure Analysis and Pretended Reflections
After inspecting chain of thoughtss (CoTs) of the results generated by gmn3, we find that during the image generation process, gmn3 performs a three-step CoT before presenting the final result: it first considers the task and requirements, then generates a draft image, and finally checks the draft against the requirements and reflects on the result. Such a process seems to imply quality control and iterative refinements. However, as we can see in Fig. 4, a low quality mask could pass its final check in its CoT.
Figure 7: More failure instances. Top Right: The palette of related labels. Bottom left: gmn3 correctly identified the left and right eyebrows while confused about the left and right eyes. Bottom right: It mislabeled the hand as cloth.
We further investigate such failures with more attempts on the same image. Although we could not reproduce the extreme failure in Fig. 4, we present two interesting instances in Fig. 7. For the partially incorrect case (bottom left in Fig. 7), gmn3 labeled the right and left eyebrows correctly but mislabeled the eyes, which are in close proximity. In its CoT, however, it explicitly concluded that “I’ve verified that the segmentation mask strictly adheres to all user-specified constraints. Facial feature delineation, including the critical left/right reversal rule, is accurate…” Such pretended quality control also occurred in the second case (bottom right in Fig. 7), in which it misclassified the hand as cloth. In its CoT, it reflected: “… The background, hair, skin, neck, nose, eyebrows, eyes, lips and cloth were all correctly segmented. I’m satisfied that this fulfills all requirements.
The reflections of the model are merely pretending, blindly affirming the correctness of the result. This may be a fundamental flaw in its multimodal reasoning. Such examples may also represent adversarial attacks that target potential flaws and bypass material checking.
4.4 More experiments on coco
Even in the failure cases of gmn3 shown in Figs. 4 and 7, most of the results are still reasonably good, which makes us wonder whether face parsing on celeb is too easy for advanced OMMs such as gmn3. Therefore, we tested the performance of gmn3 and gmn25 and compared them with 1former on coco. Compared to celeb, coco is much more challenging, since its number of classes is much larger (144 vs. 19). We present and analyze the best and worst predictions of gmn3 in this section.
Figure 8: Best prediction (Top Right) of gmn3 on coco, achieving F1 score $0.269$. The bottom row showcases other four attempts.
Figure 9: Worst prediction (Top Right) of gmn3 on coco, achieving F1 score $0.0$. The bottom row showcases other four attempts.
The best result (Fig. 8) is interesting in that gmn3 seems to give up on drawing a detailed mask; yet, the F1 score is higher than those of other attempts, which appear more plausible because, in this example, the area with the “bottle” label is dominantly larger than the areas of other labels. Additionally, the issue of pretended reflections persists in this case, as the reflection of the model is:
I’m currently verifying the semantic segmentation mask’s consistency with the input image. I’ve analyzed the color mapping for the “bottle” category, expecting a uniform light blue representation. The generated output is a solid light blue square, which accurately reflects the category distribution in the original image. There appear to be no other categories present in the output.
As seen in the reference mask and the results of other attempts, there are categories other than “bottle” in the image, and gmn3 can correctly identify some of them, such as “background”. However, its reflection fails to initiate a corrective action.
Moreover, the CoT of the worst result (Fig. 9) reveals a fundamental flaw in gmn3’s visual perception module and visual reasoning process. It mistakenly identifies the entire picture as “net”, while in other parallel attempts, the vision system can identify details like fences, which implies that its vision system is highly unstable with inconsistent performance. Furthermore, its multimodal CoT reasoning fails to correct its mistakes due to the issue of pretended reflections. During its examination of the generated mask, it concludes that:
This extreme example hints at further investigation. One cause may be that its internal reasoning module fails to incorporate the visual information of the image and its generated mask into its reasoning and reflection process while it learns the superficial form of reflection during training.
Despite the extreme failures, the outcomes of multiple attempts from gmn3 are still plausibly good, in contrast to the results of gmn25, as shown in Fig. 10, where gmn25 completely failed to generate an SS mask. Our quantitative results in Fig. 11 align with our qualitative analysis as well. For mIoU and Dice, please refer to Figs. 5 and 6 in Suppl. C.
Figure 10: Comparison between the Results of Different Models on coco. Their short code names and F1 scores are on the top of the images. We picked the image on which 1former achieved the highest F1 score $0.8096$ while gmn3 achieved $0.1146$.
@@ 4.4 More experiments on coco
Figure 11: F1 Scores for experiments on coco. For $p=[3,5]$, we ask OMMs to try 3 or 5 times and select the best result in these attempts. As 1former contain no stochastic components, we did not run more attempts. Due to their poor performance in previous experiments, we did not run experiments for other OMMs.
5 Limitations and Discussion
We benchmark recent OMMs on the subsets of CelebAMask-HQ Lee et al. (2020) and COCO Lin et al. (2015) in terms of F1, mIoU, and Dice scores. We further analyse the results and discover the limitations of existing OMMs. Here, we discuss our limitations and future directions in terms of data, models, and metrics.
Datasets: Due to resource constraints, we did not conduct the experiments on the entire CelebAMask-HQ Lee et al. (2020) and COCO Lin et al. (2015) datasets. The quantitative results may be biased towards these subsets. We did not experiment with the models on more segmentation datasets (e.g., SA-CO Carion et al. (2025)). However, PixelArena can be easily extended to other segmentation datasets. Therefore, we leave this as future work.
Data Refinement and Dataset Development: Because many of the results provided by OMMs (e.g., results by Gemini 3 Pro Image (gmn3) in Figs. 23, and 5) are good enough to serve as initial drafts for human annotation or even better than human annotations, in the future, we can use OMMs to re-examine the data and labels in existing segmentation benchmarks and refine their data quality, as well as improve the efficiency of annotation in new benchmarks.
Model Selection: We do not benchmark against RAS Cao et al. (2025), SAM Agent Carion et al. (2025), or SAM4MLLM Chen et al. (2024). The performance difference between native generation (i.e., ours) and these model integration methods remains to be seen. However, similar to specialized models (e.g., segface and 1former), these methods have unfair advantages over OMMs since they are designed and trained for specific SS tasks. Therefore, similar to the results generated by the specialized models, the results from RAS Cao et al. (2025), SAM Agent Carion et al. (2025), and SAM4MLLM Chen et al. (2024) are not comparable to those from OMMs.
Better Prompts for OMMs: As seen in Figs. 3 and  4, one subtle but noticeable difference is in the eye areas. In reference masks, the masks for the eyes cover only the eyeballs, but in the predictions of gmn3, the masks cover most of the periorbital regions. Such a difference may be due to our under-specification of the task, as we did not mention in our prompt whether it should label only the eyeballs. If we provide more detailed instructions, we may be able to improve performance further.
OMM Research: The reason why the shuffled color encodings improve the performance of gmn3 remains an interesting topic that may be closely related to its vision system and visual reasoning capabilities. Furthermore, if we could gain access to the source code and weights of gmn3, the examples of fake quality control may be valuable for mechanistic interpretability research Bereska and Gavves (2024) to recover the mechanisms of its internal visual and generative systems.
Metric Design: As we have seen in Fig. 5, the score discrepancy between segface and gmn3 is significant; however, it does not reflect the visual similarity between the two masks. As OMMs may be significantly valuable in many applications (e.g., refining the data of existing SS datasets), we need metrics that are better than F1 Score and mIoU to evaluate their performance and guide related research.
6 Conclusion
We present PixelArena, in which we propose using segmentation tasks to probe the pixel-precision visual intelligence of advanced OMMs. We use semantic segmentation tasks on CelebAMask-HQ Lee et al. (2020) and COCO Lin et al. (2015) to test the PPVI of frontier OMMs (i.e.DeepMind (2025b, a); OpenAI and et al. (2024); Cui et al. (2025); Li et al. (2025)). With our benchmark, we find that Gemini 3 Pro Image represents a major breakthrough in this front. With qualitative and quantitative results, it demonstrates superior performance under our zero-shot setting. We also present failure cases of these models and discuss their failure modes, which shed light on potential future research directions in dataset development, OMM research, and metric design.